# Zero shot VLMs for hate meme detection: Are we there yet?

### v2

@@ 8.1 Occlusion based result interpretation
CASE 4: Original meme misclassified as negative and none of the occluded versions resulted in the correct prediction (see Table 4 for some representative examples). FHM dataset: Majority of the memes contain implicit hate. Individually neither the image nor the embedded text in the memes portray anything harmful. Most text have words like ‘dishwater’, ‘sandwich maker’, ‘girl’, ‘wive’, ‘girlfriend’. The images in these memes have cheerful faces of women with no vulgarity. When both the image and text are taken together they portray hate and, quite naturally, the model has difficulty in identifying this implied semantics even when parts of the image are occluded. MAMI dataset: Once again these memes seem to bear implicit misogynistic content. Words like ‘dishwater’, ‘sandwich maker’, ‘kitchen’, and those referring to implicit body shaming appear in the embedded text. The model does not seem to have the requisite reasoning ability to infer the correct class of the meme and occlusion naturally does not come to any help. HARM P+C dataset: A large majority of memes in this group portray fake conversations amongst political leaders. These conversations are implicitly harmful and thus the model misclassifies both the original and the occluded memes.
8.2 Typology of the error cases
Figure 1: Misclassification to hateful memes in FHM dataset.
Figure 2: Misclassification to misogynistic memes in MAMI dataset.
Figure 3: Misclassification to harmful memes in HARM C+P dataset.
Figure 4: Misclassification to not hateful memes in FHM dataset.
Figure 5: Misclassification to not misogynistic memes in MAMI dataset.
Figure 6: Misclassification to not harmful memes on HARM C+P dataset.
While the previous section allowed us to obtain reasons for misclassification using the occlusion approach, it is largely manual. In this section we present an automatic method to induce the cases we observed earlier.
As a first step, for each dataset, we organise the misclassified data points into two groups – (a) misclassified as positive (hateful/misogynistic/harmful) and (b) misclassified as negative. Next for the data points for each group of each dataset we first obtain embeddings of the meme image + OCR text using the clip-ViT-B-32 model. We then run multimodal BERTopic  [^2] https://maartengr.github.io/BERTopic/getting_started/multimodal/multimodal.html on each group with number of clusters varying between 2 and 3 depending on the number of data points in the group. In the rest of this section we present the results obtained for each group.
Misclassified as positive - FHM dataset: We observe that the topics in the first cluster (Figure 1) contains the word ‘gorilla’. Nevertheless, we did not find images or induced topic words with profanity in this cluster. The second and third image clusters and the induced topic words cover most of our observations discussed in CASE 2. Some instances of CASE 1 are also observed. MAMI dataset: All three clusters (Figure 2) largely correspond to CASE 1 discussed in previous subsection. None of the clusters correspond to CASE 2 discussed previously which contained nudity/vulgarity. HARM P+C dataset: We observe in Figure 3 that the first and second image clusters contain majority of the memes with multiple images stacked together and have very long texts. This is analogous to CASE 1. Some of the topic words obtained in these clusters are ‘quarantine’, ‘coronavirus’, ‘virus’, ‘china’, ‘corona’ which possibly confuses the model predictions. The third cluster is analogous to CASE 2 and depicts images which are indeed harmful.
Misclassified as negative - FHM dataset: Interestingly, in the first cluster, we get almost the same instances that we found in CASE 3 (Figure 4) where the images have irregular shapes and very small sized embedded text. The second cluster covers the point we discussed in CASE 4 and identifies topic words like ‘sandwich’, ‘dishwater’, ‘soap’ with relevant associated images. MAMI dataset: We observe that the first cluster (Figure 5) mostly resembles the CASE 4 discussed earlier. The second and third clusters contain images and topic words corresponding to CASE 3. HARM P+C dataset: The two clusters in Figure 6 do not seem to be fully analogous to any of the cases. However, the second cluster/topic words partially resembles CASE 3.
@@ 8.2 Typology of the error cases
Overall we believe the above two subsections together provide invaluable insights into what are the systematic error patterns that VLMs are vulnerable to. These insights can be directly used in developing safety guardrails as opposed to expensive repeated fine-tuning.
9 Conclusion
@@
We present a comprehensive study of popular open source VLMs for hateful meme detection, considering eight different prompt variants. For this study, we utilize four datasets covering various hate dimensions and observe that model performance varies based on datasets and prompts used. Furthermore, we also propose an approach to select the best model and prompt combination that generalizes well over considered datasets. Finally we present a systematic method to induce a typology of the errors committed by such VLMs which could have a long-term impact on how safeguarding approaches should be built in future.
10 Limitations
@@
Our work has a few limitations. First, we conducted our experiments on English meme datasets and did not assess the model’s capability for multilingual hate meme detection. Second, although we experimented with various prompt settings to identify misclassification patterns, these prompt variants are not exhaustive, and numerous other variants could be explored. Despite this, we are confident that our range of prompts can unveil the actual performance of VLMs in hate meme detection as they cover various broad meta-aspects. Third, we did not use hate meme datasets tailored explicitly for this task by fine-tuning the VLMs. In future, we plan to address these limitations.
11 Ethical statement
Our analysis refrains from attempting to trace users involved in disseminating hate, and we do not intend to harm any individuals or target communities. All experiments were thoroughly conducted using datasets crafted from prior research. Our primary focus was to assess the efficacy of large VLMs in hate meme detection, aiming to pinpoint potential areas for future enhancement.
@@ A.2 MAMI dataset
A.3 HARM-C and HARM-P datasets
  • harmful: Multi-modal units consisting of an image and a piece of text embedded that has the potential to cause harm to an individual, an organization, a community, or the society more generally. Here, harm includes mental abuse, defamation, psycho-physiological injury, proprietary damage, emotional disturbance, and compensated public image.
  • not-harmful: Multi-modal units consisting of an image and a piece of text embedded which does not cause any harm to an individual, an organization, a community, or the society more generally.
@@ Appendix B Prompt strategies
1. ‘positive_label’ - "Definition of ‘positive_label’ corresponding to dataset"
2. ‘negative_label’ - "Definition of ‘negative_label’ corresponding to dataset"
@@
1. ‘positive_label’ - "Definition of ‘positive_label’ corresponding to dataset"
2. ‘negative_label’ - "Definition of ‘negative_label’ corresponding to dataset"
@@
1. ‘positive_label’ - "Definition of ‘positive_label’ corresponding to dataset"
2. ‘negative_label’ - "Definition of ‘negative_label’ corresponding to dataset"
@@
1. ‘positive_label’ - "Definition of ‘positive_label’ corresponding to dataset"
2. ‘negative_label’ - "Definition of ‘negative_label’ corresponding to dataset"
@@
Table 5: Representative examples for different prompt variants. Definition for corresponding labels can be picked from Appendix  A. ‘positive_label’ and ‘negative_label’ will be replaced by corresponding labels as per the dataset. Note: Important text in each prompt variant has been indicated in boldface.
@@ Appendix C Reproducibility steps
Datasets: All four datasets which we have used are commonly used for hateful/misogynistic/harmful meme detection tasks. The links to these datasets can be found here – (FHM) [^3] https://www.kaggle.com/datasets/parthplc/facebook-hateful-meme-dataset, (MAMI) [^4] https://github.com/TIBHannover/multimodal-misogyny-detection-mami-2022?tab=readme-ov-file and (HARM-C & HARM-P) [^5] https://github.com/LCS2-IIITD/MOMENTA.
@@
API and passed the image and text to the processor before feeding it to the model. Here we passed model_checkpoint corresponding to the models we used (as discussed in section 4).
Quantization: We used HuggingFace APIs for quantization [^6] https://huggingface.co/docs/transformers/en/main_classes/quantization. We specifically used bitsandbytes [^7] https://huggingface.co/docs/bitsandbytes/index library to setup 8-bit quantization.
Embedded text (OCR): We used the OCR text already present in the dataset for our experiments on memes.
@@ Appendix D Examples of ambiguous outputs
Example output for hateful meme: hateful - This meme is hateful because it makes fun of a woman’s appearance and implies that she is not capable of being a good mother.’

### v3

@@ Results
CASE 4: FHM dataset: Majority of the memes contain implicit hate. Individually neither the image nor the embedded text in the memes portray any hate. Most text have words like ‘dishwater’, ‘sandwich maker’, ‘girl’, ‘wive’, ‘girlfriend’. The images in these memes have cheerful faces of women with no vulgarity. When both the image and text are taken together they portray hate and, quite naturally, the model has difficulty in identifying this implied semantics even when parts of the image are occluded. MAMI dataset: Once again these memes seem to bear implicit misogyny. Words like ‘dishwater’, ‘sandwich maker’, ‘kitchen’, and those referring to implicit body shaming appear in the embedded text. The model does not seem to have the requisite reasoning ability to infer the correct class of the meme and occlusion naturally does not come to any help. HARM P+C dataset: Drawing conclusion from this case is tricky. However, we observe a common set of memes that portray fake conversations amongst political leaders. These conversations are implicitly harmful and are present in a satirical manner. The model misclassifies both the original and the occluded memes. BHM + HinGlish datasets: Interestingly, we observed many samples wrongly annotated for HinGlish while BHM contained stacked image based memes.
Actionable evaluation: Typology of the error cases
While the previous section allowed us to obtain reasons for misclassification using the occlusion approach, it is largely manual. In this section we present an automatic method to induce the cases we observed earlier. This investigation is motivated by the concept of actionable evaluation through error typology induction introduced in (Font Llitjós, Carbonell, and Lavie 2005; Vilar et al. 2006). As a first step, for each dataset, we organise the misclassified data points into two groups – (a) misclassified as positive (hateful/misogynistic/harmful/offensive) and (b) misclassified as negative. Next for the data points for each group of each dataset we first obtain embeddings of the meme image + OCR text using the clip-ViT-B-32 model. We then run multimodal BERTopic [^10] https://maartengr.github.io/BERTopic/getting˙started/multimodal/multimodal.html on each group and bifurcate them into two clusters. Two clusters are specifically chosen to analyze the distribution of two cases (CASE 1/CASE 2 or CASE 3/CASE 4) which we covered in the previous subsection. We present our results in Figure 2111111A similar analysis of the model generated explanations are presented in the Appendix. and in the rest of the section we attempt to make inferences from the obtained results. Note that here we present the results of English datasets; error typology of BHM and HinGlish is updated in Appendix.
(i) We observe that across each cluster (with corresponding topic words), outputs of GPT-4o fall into CASE 2 and CASE 4, which naturally follows from the rigidity analysis we have performed. We observe similar patterns for LLaVA-1.5 13B version for the FHM dataset; for the other datasets, the outputs of LLaVA-1.5 13B fall mostly into CASE 1 and CASE 3. This corroborates with our conclusion from human evaluation done earlier.
(ii) The topic words and the corresponding clusters for misclassification to hateful memes for FHM are nearly same for both models. However, for GPT-4o the topics obtained are more fine-grained compared to LLaVA-1.5 13B in the following cases – misclassification to not-hateful class for FHM dataset and misclassification to either of the classes for the other datasets.
(iii) For HARM-C + P, we observe that for keywords related to Covid-19, both models tend to misclassify to harmful. Note that this is an important observation, since we arrived at no such conclusion from the human annotations.
(iv) The implicit keywords like dishwater, sandwich, polish, and maker are instances where current VLMs fail.
(v) GPT-4o in case of misclassification to not-misogynistic for the MAMI dataset fails to understand the implicit hateful references to words like food, hot, and fat.
(vi) Both LLaVA-1.5 13B and GPT-4o in case of misclassification to hateful memes topics contain keywords related to religion. Thus non-hateful memes containing common target words can often get misclassified.
Action items
@@
Overall we believe the above two subsections together provide invaluable insights into what are the systematic error patterns that VLMs are vulnerable to. We summarize our analysis and discuss some action items:
(i) Annotation quality: Our study reveals that the annotation quality of datasets is a concerning problem. No matter how fancy and sophisticated the models are, the veracity of the results will remain questionable if one cannot ensure the robustness of annotation quality in future. Systematic phase-wise incremental pilots should be undertaken (similar to (Mathew et al. 2021)) to improve the annotation quality. At the end of each pilot, the quality of each annotation needs to be reviewed. One particularly important trend of good quality annotation is that the inter-annotator agreement should improve at the end of each pilot. Errors committed by the annotators in the early stages of the pipeline should be pointed to them so that these are not repeated in future. Suitable explanations need to be presented to the annotators so that they can understand their errors better. Continuously underperforming annotators, especially in the early pilot stages need to be removed from the pool. Finally, benchmark tasks can be evaluated against the annotated data after every pilot stage and performance can be monitored. If there is a drop in the performance, the annotations need to be revisited and corrective steps need to be taken.
(ii) Occlusion as blackbox model audit: Since social media memes include both images and embedded text, typical interpretability schemes like LIME (Ribeiro, Singh, and Guestrin 2016), SHAP (Lundberg and Lee 2017), GRADCAM (Selvaraju et al. 2017) etc. are not quite apt for them. Moreover, these methods cannot be applied on closed source models like GPT-4o. In contrast, our simple yet effective occlusion based strategy is able to reveal highly relevant interpretations of why VLMs go wrong and this approach is easily generalizable to other similar downstream tasks.
(iii) Error typology as safety guard rail: On social media platforms, the content of memes change frequently. Thus costly finetuning of the models is repeatedly required to adapt to these new changes. The meta-level categories of memes like the topic clusters that we induced are far less volatile and any newly formed hate meme can be easily mapped to one of these clusters. Therefore it should suffice to feed these clusters into the design pipeline of the safety guardrails for the VLMs thus eliminating the need for regular finetuning of these models. Further, it can be brewed to design functionality based trees and preference datasets for VLM alignment. For instance, one can attempt to prepare error type aware preference datasets for an RLHF style safety fine-tuning.
Conclusion
@@
We present a comprehensive study of popular VLMs on hateful memes, spanning eight different prompt variants. For this study, we utilize six datasets spanning three languages and covering various toxicity dimensions and observe that model performance varies based on datasets, language and prompts used. Furthermore, we also propose an approach to select the best model and prompt combination that generalizes well over considered datasets & languages. Finally we present a systematic method to induce a typology of the errors committed by such VLMs which could have a long-term impact on how safeguarding approaches should be built in future.
Limitations
@@
Our work has a few limitations. First, although we experimented with various prompt settings to identify misclassification patterns, these prompt variants are not exhaustive, and numerous other variants could be explored. Despite this, we are confident that our range of prompts can unveil the actual performance of VLMs in hate meme detection as they cover various broad meta-aspects. Second, we did not use hate meme datasets tailored explicitly for this task by fine-tuning the VLMs because of the issues pointed out in the introduction section. Finally, all our experiments are performed by manually tuning the temperature parameter which is fixed for a specific model; to perform fair evaluation across datasets and prompt variants. Other values can also be tested. In future, we plan to address these limitations.
@@ Appendix A Definitions
HARM-C and HARM-P datasets
  • harmful: Multi-modal units consisting of an image and a piece of text embedded that has the potential to cause harm to an individual, an organization, a community, or civil society more generally. Here, harm includes mental abuse, defamation, psycho-physiological injury, proprietary damage, emotional disturbance, and compensated public image.
  • not-harmful: Multi-modal units consisting of an image and a piece of text embedded that does not cause any harm to an individual, an organization, a community, or society more generally.
BHM dataset
  • hateful: If it explicitly intends to denigrate, vilify, harm, mock, abuse any entity based on their gender, race, ideology, belief, social, political, geographical and organizational status.
  • not-hateful: If it is not hateful and follows social norms. community, or the society more generally.
HinGlish dataset
  • offensive: A meme will be categorized as offensive if it either explicitly or implicitly dehumanizes, degrades, insults, or attacks any individual or group based on attributes, such as gender, nationality, sexual orientation, ethnicity, race, skin color, health condition.
  • not-offensive: A meme that is not offensive and follows social norms.
@@ Appendix B Prompt strategies
1. ‘positive_label’ - “Definition of ‘positive_label’ corresponding to dataset”
2. ‘negative_label’ - “Definition of ‘negative_label’ corresponding to dataset”
@@
1. ‘positive_label’ - “Definition of ‘positive_label’ corresponding to dataset”
2. ‘negative_label’ - “Definition of ‘negative_label’ corresponding to dataset”
@@
1. ‘positive_label’ - “Definition of ‘positive_label’ corresponding to dataset”
2. ‘negative_label’ - “Definition of ‘negative_label’ corresponding to dataset”
@@
1. ‘positive_label’ - “Definition of ‘positive_label’ corresponding to dataset”
2. ‘negative_label’ - “Definition of ‘negative_label’ corresponding to dataset”
@@
Table 8: Employed Prompts: Representative examples for different prompt variants. Definition for corresponding labels can be picked from AppendixA. ‘positive_label’ and ‘negative_label’ will be replaced by corresponding labels as per the dataset. Note: Important text in each prompt variant has been indicated in boldface.
@@ Appendix C Reproducibility steps
Datasets: All six datasets that we have used are commonly used for hateful/misogynistic/harmful/offensive meme detection tasks. The links to these datasets can be found here – (FHM) [^12] https://www.kaggle.com/datasets/parthplc/facebook-hateful-meme-dataset, (MAMI) [^13] https://github.com/TIBHannover/multimodal-misogyny-detection-mami-2022?tab=readme-ov-file, (HARM-C & HARM-P) [^14] https://github.com/LCS2-IIITD/MOMENTA, (BHM) [^15] https://github.com/eftekhar-hossain/Bengali-Hateful-Memes/blob/main/README.md, and (HinGlish) [^16] https://github.com/Gitanjali1801/CM˙MEMES.
@@
API and passed the image and text to the processor before feeding it to the model. Here we passed model_checkpoint corresponding to the models we used (as discussed in main content).
Quantization: We used HuggingFace APIs for quantization [^17] https://huggingface.co/docs/transformers/en/main˙classes/quantization. We specifically used bitsandbytes [^18] https://huggingface.co/docs/bitsandbytes/index library to setup 8-bit quantization.
Embedded text (OCR): We used the OCR text already present in the dataset for our experiments on memes. For BHM dataset, we translate Bengali OCR text to English using AI4Bharat [^19] https://ai4bharat.iitm.ac.in/ models.
BerTopic: We set random_state equal to 42 in Umap while generating clusters.
License agreement: We agreed to every license before accessing any model/dataset. Since we used open-source models using HuggingFace, we agreed to their license and for the datasets, we thoroughly read the license and used them as per the corresponding agreement. Please navigate to corresponding mentioned dataset URLs/model checkpoints to get complete license details.
@@ Appendix D Examples of ambiguous outputs
Example output for hateful meme: hateful - This meme is hateful because it makes fun of a woman’s appearance and implies that she is not capable of being a good mother.
@@
Appendix E Multimodal BERTopic image clusters
The image clusters of the error typology induction for FHM, MAMI and HARM C+P datasets are presented in Figure 3.
Figure 3: Typology Clusters: Upper panel images: GPT-4o error clusters. Lower panel images: LLaVA-1.5 13B error clusters. Each cluster and dataset are separated by dashed lines.
Appendix F BHM and HinGlish typology
Figure 4: Typology Extended: Green circles represent misclassification to positive label; red signifies misclassification to negative label. Each set of misclassification is bifurcated into two clusters. Distribution of cases, topic words and representative image cluster are shown for GPT-4o and LLaVA-1.5 13B models, for BHM and HinGlish datasets. Important keywords in each topic are marked in bold.
From Figure 4, we surprisingly observe that both models seemingly correctly classify to offensive for BHM dataset since the induced topic – f*cking, mullah, nibba, and nibbi are indeed used to propagate hateful sentiments and are prevalent within Indian subcontinent. Also, the decreased distribution of CASE 2 and CASE 4 compared to English datasets for GPT-4o signifies its weakness for multilingual data. For HinGlish, most clusters contain common pronouns used in HinGlish language. However, for both models, the second cluster corresponding to cases of misclassification to the not offensive class has keywords related to Covid-19; this again corroborates the earlier observation from the occlusion based experiments.
Therefore, the extended study on BHM and HinGlish further demonstrates the importance of our automatic typology induction.
Appendix G Annotation guidelines
The annotation process was carried out to evaluate the effect of occlusion of misclassified memes across four cases. For each case, annotators had to bifurcate every meme into systematic error cases. These cases were defined by the authors of the paper after having a broad review of all misclassification cases. Further, it was observed that each case broadly focused on specific classes. Annotators were asked to assess each of the cases alongwith occlusion results to help us better understand the specific regions of the meme important for the models.
Four separate files were created, each corresponding to one of these cases, to document these instances and understand the nuances of model errors. Each page of the file contains the image ID, the original image, its ground truth label, the occluded images that changed the original prediction, and the corresponding predictions after occlusion. The evaluation process identified and categorized the misclassified memes into nine distinct error situations based on their visual and textual characteristics:
  1. 1.
    NHM (Not hateful memes): Memes featuring common target words (e.g., “Islam,” “migrants,” “white”), political figures, or profane language.
  2. 2.
    WA (Wrong annotation): Memes incorrectly labeled in the dataset.
  3. 3.
    SI (Stacked images): Memes with overlaid or stacked visual elements.
  4. 4.
    MCVG (Multiple color variations or grayscale): Memes containing varied color schemes, grayscale visuals, or multiple objects.
  5. 5.
    URET (Unreadable embedded text): Memes with lengthy or small-font embedded text that is difficult to read.
  6. 6.
    IM (Implicit hate meme): Memes that convey hate or harmful content implicitly, where the harmful intent is not straightforward to understand.
  7. 7.
    PFW (Perturbed or animated faces of women): Memes featuring distorted, exaggerated, or animated female faces.
  8. 8.
    NV (Nudity or vulgarity): Memes containing explicitly vulgar/nude visuals or embedded text.
  9. 9.
    FC (Fake conversation): Memes designed as fabricated dialogues or text exchanges.
The results of the annotation process, are summarized in the main content of the paper and we have provided detailed insights into the distribution of these situations across the four cases for both GPT-4o andLLaVA-1.5 13B models. For instance:
  • CASE 1: Prominently featured NHM, WA, PFW, and SI, indicating these categories were sensitive to occlusion.
  • CASE 2: Largely influenced by WA and NHM, indicating wrong annotation in datasets and stereotyping common target words.
  • CASE 3: Showed issues with NV and MCVG, where models struggled with explicit content or complex visual layouts.
  • CASE 4: Highlighted errors in processing FC and IM, emphasizing limitations in understanding fabricated contexts or subtle cues.
The annotation process sheds light on the limitations of VLMs when handling nuanced or visually complex memes. By categorizing errors and analyzing the impact of occlusion, the study underscores the importance of improving feature extraction and capturing subtle contextual cues to enhance model robustness and predictive accuracy.
Appendix H Takeaways from prompt variants
While we have covered extensive experiment and analysis in the main content, we summarize key observations here as well.
(i) Open source models tend to perform much better when vanilla as output is prompted. They seemingly confuse with the generation of explanation.
(ii) OCR text alone helps IDEFICS and the incorporation of definition generally degrades its results.
(iii) OCR + definition works best for LLaVA-1.5 models. Surprisingly, 7B version works best with explanation as output while 13B works best with vanilla as output.
(iv) We can safely say that although open source VLMs might be good at generation, their performance as predictors needs improvement. Our study proves that although open source VLMs are good generators, a long path awaits before they match with likes of GPT-4o.
(v) We also observed that the models which generally perform better across all datasets are also more stable; hence GPT-4o had the least standard deviation across all models considered.
(vi) On multilingual datasets, even GPT-4o struggles in terms of performance.
(vii) InstructBLIP models generate very large outputs which leads to ambiguity. This behaviour considerably increases in case of explanation as output.
Appendix I Examples of wrong annotation
Figure 5 presents fifteen memes; nine from MAMI and six from FHM. All these memes are marked as either misogynistic or hateful by human annotators employed in our work. All these fall in CASE 2 i.e., none of the occluded samples were able to correct the prediction. We can observe that GPT-4o does not only predict these cases correctly, but also provides reasonable justification.
Figure 5: Examples of Wrong Annotation: Fifteen examples, nine from MAMI and six from FHM are shown for CASE 2 of GPT-4o with def + OCR as input and explanation as output. Output of the model is also provided.
Appendix J Analysis of GPT-4o explanations
Figure 6: Typology Explanation: Green circles represent misclassification to positive label; red signifies misclassification to negative label. Each set of misclassification is bifurcated into two clusters. Distribution of cases and topic words are shown for GPT-4o for FHM, MAMI, HARM-C + P, BHM and HinGlish datasets. Important keywords in each topic are marked in bold.
In Figure 6, we provide the induced typology of GPT-4o explanations for misclassified samples. We use the default BERTopic pipeline to generate typology alongwith clusters and summarize our key observations:
(i) The induced topics are not very diverse compared to what we obtained previously (refer to Figures 2 and 4). Although these topics provide insights into the generation capability of the model, they cannot be effectively brewed as safety guardrails.
(ii) From the misclassification to negative labels (red circles in Figure 6), we observe that models generate words like ‘humorously’ and ‘humorous’. Drawing from occlusion based analysis we conclude that models generally consider implicitly hateful memes as humorous and are unable to identify hidden hateful elements.
(iii) Further we also observe from the misclassification to negative labels that the model is seemingly over aggressive to classify memes targeting ‘group’, ‘society’ and ‘organization’ as not-(hateful/harmful/offensive).
(iv) The topic ‘stereotype’ is only present in misclassification to positive label for MAMI, BHM and HinGlish datasets. Interestingly, ‘religion’ and ‘religious’ also fall under misclassification to positive label for FHM & BHM (both datasets contain hateful memes). This reinforces the observations from our occlusion based study whereby the models seem to excessively stereotype religion and forcefully misclassify.
(v) Another interesting observation is the presence of words like ‘implies’ and ‘implying’ only in the topic clusters for miscalssification to positive labels – this means that models like GPT-4o attempt to stress on defending its classification to positive labels.
(vi) GPT-4o when given definition as input generates relevant explanations, containing the words like ‘dehumanizing’, ‘objectification’, ‘stereotypes’, ‘shaming’ and ‘defamation’ among many others. This tells us that after analyzing the image and OCR text, GPT-4o is able to lucidly apply reasoning to generate task specific and relevant output.
(vii) Surprisingly, the case distribution of clusters highly align with what we concluded in actionable evaluation study. This concludes that the rigidness distribution is just not limited to inputs but is propagated to outputs as well.